OpenStudy (anonymous):

What is the value of x? 4(6 + 7) = 4 • 6 + x • 7

OpenStudy (anonymous):

fan+medal

OpenStudy (itsbribro):

x=4

OpenStudy (itsbribro):

did you try PEMDAS

OpenStudy (anonymous):

doing that now

OpenStudy (itsbribro):

4(6+7)=4*6+x*7 4(13)=4*6+x*7 52=4*6+x*7 52=24+x*7 52=24+4*7 52=24+28 52=52

OpenStudy (itsbribro):

if i messed up anywhere sorry

OpenStudy (anonymous):

so what i have so far is 52=(24)+x*7

OpenStudy (itsbribro):

ok so your next step is you take 52-24=

OpenStudy (itsbribro):

=28 so 7*? =28

OpenStudy (anonymous):

28

OpenStudy (itsbribro):

ok so 7*?=28

OpenStudy (anonymous):

7*4=28 ?

OpenStudy (itsbribro):

yes so you simplify that and get 52=24+28

OpenStudy (itsbribro):

so waht is 24+28?

OpenStudy (anonymous):

52

OpenStudy (itsbribro):

good now 52=52 you automatically know that x=7 because x*4=28 and 28+24=52

OpenStudy (itsbribro):

if that makes any sense

OpenStudy (anonymous):

lol soo x=7 ??

OpenStudy (itsbribro):

i meant 4 sorry

OpenStudy (itsbribro):

lol x=4 sorry

OpenStudy (anonymous):

I have another one.

OpenStudy (anonymous):

Which expressions show how 5 • 37 can be rewritten using the distributive property? Choose all answers that are correct. A. 5 • 30 + 7 B. 5 • 30 + 5 • 7 C. 5 • 40 – 5 • 7 D. 5 • 40 – 5 • 3

OpenStudy (itsbribro):

idek but i think its A and D

OpenStudy (anonymous):

Simplify: 6m + 7n + 5m – 3n